OpenStudy (anonymous):

A parabola has the equation –5x^2 + 30x – y – 34 = 0. Which equation is that of the parabola in vertex form?

OpenStudy (mertsj):

\[y + 34=-5(x^2-6x)\]

OpenStudy (mertsj):

\[y+34 -45 = -5(x^2 - 6x +9)\]

OpenStudy (mertsj):

\[y=-5(x-3)^2+11\]